# Dark Matter (DM); Dark Stars (DSs); High-z galaxies; Supermassive Black Holes (SMBHs), Little Red Dots (LRDs); Blue Monsters; Over-massive BH galaxies (OBGs)

### v1

@@ Abstract
[…]
@@
Dark Matter (DM); Dark Stars (DSs); High-z galaxies; Supermassive Black Holes (SMBHs), Little Red Dots (LRDs); Blue Monsters; Over-massive BH galaxies (OBGs)
@@ 1 Introduction
JWST uncovered a larger than expected number of high redshift Lyman break galaxies which are very bright yet extremely compact ($r_{eff}\lesssim 400$ pc) (e.g. JWSTBLueMonsters:2025A&A...694A.286F, and references therein). If interpreted as galaxies, those “Blue Monsters” BlueMonsters:2023MNRAS.520.2445Z are too large for their epoch ($M_{\star}\gtrsim 10^{9}M_{\odot}$) and their large comoving number density (BlueMonstersNumbers:2024ApJ...965...98C; BlueMonstersHighNumber:2025ApJ) is very hard to explain with standard, pre-JWST models of the formation of the first stars and assembly of the first galaxies. While there are uncertainties, such as the estimates of how massive single stars are (Steinhardt:2023Templates; Lapi:2024tgm), or accurately accounting for the presence of interstellar dust (Ferrara:2022), if this tension persists as JWST collects more and more data it would be a clear indication that our current understanding of the formation and evolution of the first stars and galaxies is, at best, incomplete.
@@
The origin of the Supermassive Black Holes (SMBHs) powering the most distant observed Quasars (QSOs) remains an open question, deepened by JWST data. The most striking such example is UHZ1, the most distant quasar observed ($z\sim 10$). JWST and Chandra data shows that UHZ1 harbors an SMBH with $M_{BH}\sim 10^{7}M_{\odot}$. Such enormous Black Holes (BHs) should not have existed so early if they were seeded by regular, nuclear powered stars (Bogdan:2023UHZ1). UHZ1 is just the tip of the iceberg when it comes to the distant quasars observed, each harboring SMBHs that should not have had time to grow as massive as they are (e.g. Wang:2019; Inayoshi:2020; Lupi:2021). If one insists on a regular nuclear powered stellar seed then long period, sustained super-Eddington accretion rates are required to explain the SMBHs powering UHZ1 and the other most distant QSOs. More likely, the data implies the need for heavy Black Hole seeds, at high redshifts (Bogdan:2023UHZ1). Direct Collapse Black Holes (DCBHs) (e.g. Loeb:1994wv; Belgman:2006; Lodato:2006hw; Natarajan:2017; barrow:2018; Whalen:2020; Inayoshi:2020) and Dark Stars (e.g. Spolyar:2008dark; Freese:2008ds; Freese:2008cap; Iocco:2008cap; Freese:2010smds)  [^1] In the literature sometimes Dark Stars are called Pop III.1 stars (e.g. Banik:2019; TanPopIIISMDS2025; PopIII1SMBHs:2024; PopIII1SMBSs:2019) naturally generate such heavy seeds.
@@
Lastly, JWST uncovered a completely new class of objects at the cosmic dawn era, the so called little red dots (LRDs) (e.g. LDRs:2024ApJ...963..129M; LRDsCensus:2024ApJ...968...38K; LRDs:2025ApJ...991...37A, and references therein). They are extremely compact, with effective radii $\lesssim 200$ pc, and thus unlikely to be galaxies, as that would require an impossibly dense packing of stars. While there is no consensus yet as to what the LRDs are, several key, puzzling facts are known: they emit weakly in UV, and not at all in X-rays; this cannot be explained by dust obscuration, in view of the lack of re-emissions at higher wavelengths, when LDRs are observed with MIRI or ALMA (LDRsNoALMA:2025ApJ...990L..61C). Thus, it is very unlikely that LRDs are either galaxies or Quasars. The picture that seems to emerge is that many LRDs consists of a dense gas envelope that encircle an active BH; the black hole’s emitted radiation provides energy that supports this gas layer, which, in turn, has sufficient density to capture and absorb high-energy ultraviolet light and X-ray radiation (e.g. LRDsBHStar:2025arXiv250316596N; LRDBHStar:2025A&A...701A.168D; LRDsBHStars:2025arXiv250709085B). In the literature, the formation of such objects (also known as quasi-stars) has been hypothesized, in the context of the Direct Collapse Black Hole scenario, as one of the stages of the formation of SMBHs seeded by the collapse of nuclear powered Supermassive Stars (Belgman:2006). On the other hand, SMBHs seeded by the collapse of SMDSs could, in principle lead to a very similar configuration, as the macro scale astrophysics is very similar: supermassive stars that eventually run out of fuel (nuclear for DCBHs vs DM for SMDS) or attain a sufficiently large mass to become unstable under general relativistic instabilities. When either of those two conditions is met, the supermassive star collapses and forms a SMBH, still surrounded by the remnant of the outer layers of the star and the gas that had not yet been accreted into it prior to collapse.
@@
When combined, the three astronomical puzzles discussed above indicate that we might be on the verge of a revolution of our understanding of the cosmic dawn era. In this paper we argue that Dark Stars (DSs) are a natural solution to all three of those tensions between data and theoretical expectations. Specifically, in Sec. 2, we briefly review theoretical aspects of Dark Stars, and their observational status in view of the candidates found by Ilie:2023JADES and ilie2025spectroscopicsupermassivedarkstar. In Sec. 3 we discuss how the SMDSs candidates already identified could explain some of the “Blue Monsters” observed with JWST. In Sec. 4 we demonstrate that, in addition to DCBHs, Dark Stars provide an efficient seeding mechanism for the overly massive BHs observed at high-z, and in Sec. 5, we show how Dark Star seeded BHs could have grown to power the most distant quasars observed, such as UHZ1, J0313-1806, J1342+0928, and J1007+2115. In Sec. 6 we conclude and further elaborate how we envision Dark Stars could be part of the mystery posed by LRDs. In order to keep the main body of the paper as concise and accessible as possible we relegate most of the technical details to the following appendices: Appendix A (a brief overview of the evidence for the existence of Dark Matter), Appendix B (a supplement for our Sec. 2 on Dark Stars), Appendix C, (were we expand upon the discussion of Sec. 4 regarding DCBHs and/or DSs as SMBHs seeds), and Appendix D (where we expand upon the discussion presented in Sec. 5 regarding the mechanisms via which SMDSs or their SMBH remnants can become embedded in high-z galaxies).
@@ 2 Dark Stars (DSs): from theory to first observational hints
The first stars in the universe formed more than thirteen Gyrs ago, when molecular clouds comprised of hydrogen and helium began to gravitationally collapse at the centers of DM halos (e.g. Abel:2001; Barkana:2000; Bromm:2003; Yoshida:2006; OShea:2007; Yoshida:2008; Bromm:2009). The balance between all heating and all cooling mechanisms determines the point when the runaway collapse stops, and a proto-star in hydrostatic equilibrium forms. The very high DM densities, along with the poor cooling available in those pristine gas clouds, are the main reasons why DM burning stars (i.e. Dark Stars) could form (Spolyar:2008dark). The conditions identified in Spolyar:2008dark for the formation of DSs are expected to be met only during a relatively narrow window of redshifts, e.g. $z\in[30-10]$; as such, Dark Stars should be exceedingly rare when the Universe’s age is $\gtrsim 500$ Myrs, a fact confirmed by Ilie:2012, using Hubbbke Space Telescope (HST) data. While they are composed mainly of primordial H and He (from Big Bang Nucleosynthesis), with less than $1\%$ of their mass carried by DM particles, DSs are powered by Dark Matter heating alone.
@@
Dark Stars are expected to form at the centers of micro-DM halos, with $\sim 1$ per halo. The DM reservoir can be supplied in one of two ways. First, when the Dark Star forms, it will rely on the DM spike at the center of the micro-halo, where the progenitor DM cloud collapses. In the process, the DM densities are further enhanced, as the infall of baryons steepens the gravitational potential well, and DM orbits respond by compressing. This process can be modeled quite well via the so called Adiabatic Contraction (AC) prescription (e.g. Blumenthal:1985; Freese:2008dmdens; Gnedin:2004; Gnedin:2011). The AC mechanism relies on gravity to pull in more DM inside the star. For DM halos that are triaxial (i.e. slightly asymmetric), the AC DM reservoir can be replenished efficiently by the large population of chaotic, centrophilic DM orbits present in such halos (e.g. Valluri:2010Triaxial). Dark Stars powered by this extended Adiabatically Contracted (AC) DM reservoir can grow (via accretion) as massive as a million suns, and their temperatures stay below $\sim 20,000$ K (Freese:2010smds)  [^2] If one considers 100 GeV WIMPs annihilations as their power source. For lighter/heavier mass WIMPs SMDSs would be cooler/hotter.. Conversely, if the initial DM reservoir inside of the star is not replenished efficiently by chaotic/centrophilic orbits, then for accretion rates $\lesssim 0.01M_{\odot}/$yr the DS would have exhausted this initial AC DM reservoir by the time it reaches $\lesssim 1000M_{\odot}$, after which it enters a contraction phase. Under the assumption of non-zero interactions between DM and protons, the contracted DS can begin to efficiently capture DM from its surroundings, which in turn can provide an additional heat source. DS powered via this DM capture mechanism can also grow to become supermassive, although they will be slightly hotter than those formed via the extended AC mechanism.
@@
The equilibrium structure of DSs can be well approximated using polytropes of variable inde, ranging from $n=1.5$ (fully convective when the star is born, i.e. $M\sim 1M_{\odot}$) to $n=3$ (radiation pressure dominated, when $M_{SMDS}\gtrsim 100M_{\odot}$(Freese:2008ds; Freese:2008dmdens). Ref. Rindler-Daller:2014uja used the stellar evolution code MESA to obtain detailed models (that essentially confirmed the results using polytropes). Observable properties of dark stars cane computed from their spectra, which we modeled with TLUSTY (hubeny2017briefintroductoryguidetlusty) based on relevant macroscopic properties (i.e. Mass, Radius, composition of the stellar atmosphere).
@@
While cosmological simulations have not yet been able to reproduce the Dark Star phase, they show that DM annihilations lead to a suppression of the fragmentation of the gas clouds from which the first stars form. For instance, Refs. Ripamonti:2009xw; Smith:2012; Stacy:2014 follow the the collapse of a protostellar gas cloud, while including the possible effects of $100$ GeV annihilating WIMPs. Those studies find that collapse continues up the highest, resolution limited, baryonic number densities possible in their simulations: $n_{B}\sim 10^{14}\mathrm{cm}^{-3}$. However, for a $100$ GeV WIMP, a proto Dark Star, in hydrostatic and thermal equilibrium, is expected to form whenever $n_{B}\sim 10^{17}\mathrm{cm}^{-3}$ (Spolyar:2008dark). Therefore, contrary to claims in the literature, those studies in fact do not show that Dark Stars cannot form. However, some of those simulations find that DM annihilations can lead to the suppression of fragmentation of the gas cloud (e.g. Smith:2012; Stacy:2014), thus removing one possible obstacle to the formation of DSs (Gondolo:2013note).
@@
Dark Stars are very puffy ($R_{DS}\sim 10$ A.U.); they are also relatively cool, with temperatures ranging between a few thousands and a few tens of thousands K, depending on the DM particle mass and the formation mechanism (Freese:2010smds). This is the main reason DSs can continue to grow almost indefinitely, as long as there is a supply of baryons to accrete from and DM to annihilate. Radiative feedback effects that would shut off accretion, and which limit the mass of nuclear burning stars to a few hundred times the mas of the Sun ($M_{\odot}$), are almost irrelevant for dark stars. As such, DSs convert gas into stellar material at almost $100\%$ efficiency. This is in stark contrast with the much lower, experimentally observed, efficiency of conversion in “local” galaxies (e.g. Leroy:2008SFR; 2018Starburst). Recent JWST data indicates that many of the most distant “galaxies” ever observed must have converted gas into stars at an incredible $\gtrsim 60\%$ rate (Boylan-Kolchin:2023). Dark Stars, as well as the supermassive stars precursors to DCBHs are the only two known mechanisms via which gas is converted into stars so efficiently.

### v2

# Dark Matter (DM); Dark Stars (DSs); high-z galaxies; supermassive black holes (SMBHs); little red dots (LRDs); blue monsters; over-massive BH galaxies (OBGs)

@@ Abstract
The James Webb Space Telescope (JWST) has begun to revolutionize our view of the Cosmos. The discovery of Blue Monsters (i.e., ultra-compact yet very bright high-z galaxies) and the Little Red Dots (i.e., very compact dustless strong Balmer break cosmic dawn sources) pose significant challenges to pre-JWST era models of the assembly of first stars and galaxies. In addition, JWST data further strengthen the problem posed by the origin of the supermassive black holes that power the most distant quasars observed. Stars powered by Dark Matter annihilation (i.e., Dark Stars) can form out of primordial gas clouds during the cosmic dawn era and subsequently might grow via accretion and become supermassive. In this paper we argue that Supermassive Dark Stars (SMDSs) offer natural solutions to the three puzzles mentioned above.
@@
Dark Matter (DM); Dark Stars (DSs); high-z galaxies; supermassive black holes (SMBHs); little red dots (LRDs); blue monsters; over-massive BH galaxies (OBGs)
@@ 1 Introduction
JWST uncovered a larger than expected number of high redshift Lyman break galaxies which are very bright yet extremely compact ($r_{eff}\lesssim 400$ pc) (e.g. Ferrara et al., 2025, and references therein). If interpreted as galaxies, those “Blue Monsters” Ziparo et al. (2023) are too large for their epoch ($M_{\star}\gtrsim 10^{9}M_{\odot}$) and their large comoving number density (Casey et al., 2024; Harvey et al., 2025) is very hard to explain with standard, pre-JWST models of the formation of the first stars and assembly of the first galaxies. While there are uncertainties, such as the estimates of how massive single stars are (Steinhardt et al., 2023; Lapi et al., 2024), or accurately accounting for the presence of interstellar dust (Ferrara et al., 2023), if this tension persists as JWST collects more and more data it would be a clear indication that our current understanding of the formation and evolution of the first stars and galaxies is, at best, incomplete.
@@
The origin of the Supermassive Black Holes (SMBHs) powering the most distant observed Quasars (QSOs) remains an open question, deepened by JWST data. The most striking such example is UHZ1, the most distant quasar observed ($z\sim 10$). JWST and Chandra data shows that UHZ1 harbors an SMBH with $M_{BH}\sim 10^{7}M_{\odot}$. Such enormous Black Holes (BHs) should not have existed so early if they were seeded by regular, nuclear powered stars (Bogdán et al., 2024). UHZ1 is just the tip of the iceberg when it comes to the distant quasars observed, each harboring SMBHs that should not have had time to grow as massive as they are (e.g. Wang et al., 2019; Inayoshi et al., 2020; Lupi et al., 2021). If one insists on a regular nuclear powered stellar seed then long period, sustained super-Eddington accretion rates are required to explain the SMBHs powering UHZ1 and the other most distant QSOs. More likely, the data implies the need for heavy Black Hole seeds, at high redshifts (Bogdán et al., 2024). Direct Collapse Black Holes (DCBHs) (e.g. Loeb and Rasio, 1994; Begelman et al., 2006; Lodato and Natarajan, 2006; Natarajan et al., 2017; Barrow et al., 2018; Whalen et al., 2020; Inayoshi et al., 2020) and Dark Stars (e.g. Spolyar et al., 2008; Freese et al., 2008a, b; Iocco, 2008; Freese et al., 2010)  [^1] In the literature sometimes Dark Stars are called Pop III.1 stars (e.g. Banik et al., 2019; Sanati et al., 2025; Cammelli et al., 2024) naturally generate such heavy seeds.
@@
Lastly, JWST uncovered a completely new class of objects at the cosmic dawn era, the so called little red dots (LRDs) (e.g. Matthee et al., 2024; Kokorev et al., 2024; Akins et al., 2025, and references therein). They are extremely compact, with effective radii $\lesssim 200$ pc, and thus unlikely to be galaxies, as that would require an impossibly dense packing of stars. While there is no consensus yet as to what the LRDs are, several key, puzzling facts are known: they emit weakly in UV, and not at all in X-rays; this cannot be explained by dust obscuration, in view of the lack of re-emissions at higher wavelengths, when LDRs are observed with MIRI or ALMA (Casey et al., 2025). Thus, it is very unlikely that LRDs are either galaxies or Quasars. The picture that seems to emerge is that many LRDs consists of a dense gas envelope that encircle an active BH; the black hole’s emitted radiation provides energy that supports this gas layer, which, in turn, has sufficient density to capture and absorb high-energy ultraviolet light and X-ray radiation (e.g. Naidu et al., 2025; de Graaff et al., 2025; Begelman and Dexter, 2025). In the literature, the formation of such objects (also known as quasi-stars) has been hypothesized, in the context of the Direct Collapse Black Hole scenario, as one of the stages of the formation of SMBHs seeded by the collapse of nuclear powered Supermassive Stars (Begelman et al., 2006). On the other hand, SMBHs seeded by the collapse of SMDSs could, in principle lead to a very similar configuration, as the macro scale astrophysics is very similar: supermassive stars that eventually run out of fuel (nuclear for DCBHs vs DM for SMDS) or attain a sufficiently large mass to become unstable under general relativistic instabilities. When either of those two conditions is met, the supermassive star collapses and forms a SMBH, still surrounded by the remnant of the outer layers of the star and the gas that had not yet been accreted into it prior to collapse.
@@
When combined, the three astronomical puzzles discussed above indicate that we might be on the verge of a revolution of our understanding of the cosmic dawn era. In this paper we argue that Dark Stars (DSs) are a natural solution to all three of those tensions between data and theoretical expectations. Specifically, in Sec. 2, we briefly review theoretical aspects of Dark Stars, and their observational status in view of the candidates found by Ilie et al. (2023) and Ilie et al. (2025). In Sec. 3 we discuss how the SMDSs candidates already identified could explain some of the “Blue Monsters” observed with JWST. In Sec. 4 we demonstrate that, in addition to DCBHs, Dark Stars provide an efficient seeding mechanism for the overly massive BHs observed at high-z, and in Sec. 5, we show how Dark Star seeded BHs could have grown to power the most distant quasars observed, such as UHZ1, J0313-1806, J1342+0928, and J1007+2115. In Sec. 6 we conclude and further elaborate how we envision Dark Stars could be part of the mystery posed by LRDs. In order to keep the main body of the paper as concise and accessible as possible we relegate most of the technical details to the following appendices: Appendix A (a brief overview of the evidence for the existence of Dark Matter), Appendix B (a supplement for our Sec. 2 on Dark Stars), Appendix C, (were we expand upon the discussion of Sec. 4 regarding DCBHs and/or DSs as SMBHs seeds), and Appendix D (where we expand upon the discussion presented in Sec. 5 regarding the mechanisms via which SMDSs or their SMBH remnants can become embedded in high-z galaxies).
@@ 2 Dark Stars (DSs): from theory to first observational hints
The first stars in the universe formed more than thirteen Gyrs ago, when molecular clouds comprised of hydrogen and helium began to gravitationally collapse at the centers of DM halos (e.g. Abel et al., 2002; Barkana and Loeb, 2001; Bromm and Larson, 2004; Yoshida et al., 2006; O’Shea and Norman, 2007; Yoshida et al., 2008; Bromm et al., 2009). The balance between all heating and all cooling mechanisms determines the point when the runaway collapse stops, and a proto-star in hydrostatic equilibrium forms. The very high DM densities, along with the poor cooling available in those pristine gas clouds, are the main reasons why DM burning stars (i.e. Dark Stars) could form (Spolyar et al., 2008). The conditions identified in Spolyar et al. (2008) for the formation of DSs are expected to be met only during a relatively narrow window of redshifts, e.g. $z\in[30-10]$; as such, Dark Stars should be exceedingly rare when the Universe’s age is $\gtrsim 500$ Myrs, a fact confirmed by Ilie et al. (2012), using Hubbbke Space Telescope (HST) data. While they are composed mainly of primordial H and He (from Big Bang Nucleosynthesis), with less than $1\%$ of their mass carried by DM particles, DSs are powered by Dark Matter heating alone.
@@
Dark Stars are expected to form at the centers of micro-DM halos, with $\sim 1$ per halo. The DM reservoir can be supplied in one of two ways. First, when the Dark Star forms, it will rely on the DM spike at the center of the micro-halo, where the progenitor DM cloud collapses. In the process, the DM densities are further enhanced, as the infall of baryons steepens the gravitational potential well, and DM orbits respond by compressing. This process can be modeled quite well via the so called Adiabatic Contraction (AC) prescription (e.g. Blumenthal et al., 1986; Freese et al., 2009; Gnedin et al., 2004, 2011). The AC mechanism relies on gravity to pull in more DM inside the star. For DM halos that are triaxial (i.e. slightly asymmetric), the AC DM reservoir can be replenished efficiently by the large population of chaotic, centrophilic DM orbits present in such halos (e.g. Valluri et al., 2010). Dark Stars powered by this extended Adiabatically Contracted (AC) DM reservoir can grow (via accretion) as massive as a million suns, and their temperatures stay below $\sim 20,000$ K (Freese et al., 2010)  [^2] If one considers 100 GeV WIMPs annihilations as their power source. For lighter/heavier mass WIMPs SMDSs would be cooler/hotter.. Conversely, if the initial DM reservoir inside of the star is not replenished efficiently by chaotic/centrophilic orbits, then for accretion rates $\lesssim 0.01M_{\odot}/$yr the DS would have exhausted this initial AC DM reservoir by the time it reaches $\lesssim 1000M_{\odot}$, after which it enters a contraction phase. Under the assumption of non-zero interactions between DM and protons, the contracted DS can begin to efficiently capture DM from its surroundings, which in turn can provide an additional heat source. DS powered via this DM capture mechanism can also grow to become supermassive, although they will be slightly hotter than those formed via the extended AC mechanism.
@@
The equilibrium structure of DSs can be well approximated using polytropes of variable index, ranging from $n=1.5$ (fully convective when the star is born, i.e. $M\sim 1M_{\odot}$) to $n=3$ (radiation pressure dominated, when $M_{SMDS}\gtrsim 100M_{\odot}$(Freese et al., 2008a, 2009). Ref. Rindler-Daller et al. (2015) used the stellar evolution code MESA to obtain detailed models (that essentially confirmed the results using polytropes). Observable properties of dark stars can be computed from their spectra, which we modeled with TLUSTY (Hubeny and Lanz, 2017) based on relevant macroscopic properties (i.e. Mass, Radius, composition of the stellar atmosphere).
@@
While cosmological simulations have not yet been able to reproduce the Dark Star phase, they show that DM annihilations lead to a suppression of the fragmentation of the gas clouds from which the first stars form. For instance, Refs. Ripamonti et al. (2009); Smith et al. (2012); Stacy et al. (2014) follow the collapse of a protostellar gas cloud, while including the possible effects of $100$ GeV annihilating WIMPs. Those studies find that collapse continues up the highest, resolution limited, baryonic number densities possible in their simulations: $n_{B}\sim 10^{14}\mathrm{cm}^{-3}$. However, for a $100$ GeV WIMP, a proto Dark Star, in hydrostatic and thermal equilibrium, is expected to form whenever $n_{B}\sim 10^{17}\mathrm{cm}^{-3}$ (Spolyar et al., 2008). Therefore, contrary to claims in the literature, those studies in fact do not show that Dark Stars cannot form. However, some of those simulations find that DM annihilations can lead to the suppression of fragmentation of the gas cloud (e.g. Smith et al., 2012; Stacy et al., 2014), thus removing one possible obstacle to the formation of DSs (Gondolo et al., 2013).
While originally proposed as zero metallicity stars powered by WIMPs (Spolyar et al., 2008), Dark Stars could in principle be powered by a large variety of Dark Matter particles, as long as there is a mechanism to convert them to Standard Model particles efficiently. Out of the leading alternative models, Axions are unlikely to power Dark Stars. However, as shown by Wu et al. (2022), Self Interacting Dark Matter (SIDM) can also lead to formation of Dark Stars, with very similar properties to those powered by WIMPs. Ongoing preliminary work on number changing DM models (e.g. SIMPs, Co-SIMPs) shows promising results in terms of the potential for those kind of particles to lead to the formation of Dark Stars. In what follows we will restrict our attention (for simplicity) to DSs powered by 100 GeV WIMPs.
@@
Dark Stars are very puffy ($R_{DS}\sim 10$ A.U.); they are also relatively cool, with temperatures ranging between a few thousands and a few tens of thousands K, depending on the DM particle mass and the formation mechanism (Freese et al., 2010). This is the main reason DSs can continue to grow almost indefinitely, as long as there is a supply of baryons to accrete from and DM to annihilate. Radiative feedback effects that would shut off accretion, and which limit the mass of nuclear burning stars to a few hundred times the mas of the Sun ($M_{\odot}$), are almost irrelevant for dark stars. As such, DSs convert gas into stellar material at almost $100\%$ efficiency. This is in stark contrast with the much lower, experimentally observed, efficiency of conversion in “local” galaxies (e.g. Leroy et al., 2008; Rowan-Robinson et al., 2018). Recent JWST data indicates that many of the most distant “galaxies” ever observed must have converted gas into stars at an incredible $\gtrsim 60\%$ rate (Boylan-Kolchin, 2023). Dark Stars, as well as the supermassive stars precursors to DCBHs are the only two known mechanisms via which gas is converted into stars so efficiently.